OpenStudy (anonymous):

Vector Question: Pleeeeeease Help! If P is (-3, 2) and Q is (1, -1), find: a. Vector PQ b. magnitude or length of Vector PQ I know the answers but please tell me how to get there.

hartnn (hartnn):

\(PQ =Q(x,y)-P(x,y)= (1,-1)-(-3,2)=.....?\)

OpenStudy (anonymous):

why do u need to subtract i thought it was addition for P and Q that's where i am getting stuck

hartnn (hartnn):

because its like finding distance,  how could u have found 'p', by adding or subtracting co-ordinates ?